a company uses two types of raw material, 1 and 2 for its product. If it uses x units of 1 and y units of 2, it can produce U units of finished items, where U(x,y)=8xy+32x+40y-〖4x〗^2-〖6y〗^2. Each units of 1 costs RM 10 and each units of 2 costs RM 4. Each unit of product can be sold for RM 40. How can the company maximize its profits?
1
Expert's answer
2019-04-05T13:09:18-0400
"U(x,y)=8xy+32x+40y-4x^2-6y^2"

"R(x, y)=40(8xy+32x+40y-4x^2-6y^2)"

"C(x, y)=10x+4y"


"P(x,y)=R(x, y)-C(x, y)"

"P(x, y)=40(8xy+32x+40y-4x^2-6y^2)-10x-4y"


"P_x=320y+1280-320x-10"


"P_y=320x+1600-480y-4"

Find the critical point(s)


"P_x=0, P_y=0"

"320y+1280-320x-10=0""320x+1600-480y-4=0"

"x=y+{127\over 32}""160y=2866"

"x={3501\over 160}\approx22"

"y={2866\over 160}\approx18"

"P_{xx}=-320\lt 0""P_{yy}=-480"

"P_{xy}=P_{yx}=320"

"D=\begin{vmatrix}
 -320 & 320 \\
 320 & -480
\end{vmatrix}=51200\gt 0"

"D\gt 0, P_{xx}\lt 0". Hence, "P({3501\over 160}, {2866\over 160})" is a local maximum. Since P(x, y) has the only extrema, then "P({3501\over 160}, {2866\over 160})" is the absolute maximum.


The company may use 22 units of 1 and 18 units of 2 to maximize its profit.


"P(22, 18)=RM \ 28188."